OpenStudy (anonymous):

A triangular traffic island has a base half as long as it's height. Find the base and the height if the island has an area of 64m^2

OpenStudy (anonymous):

Please last question ppl just need help

OpenStudy (nurali):

a triangular traffic Island has a base half as long as its height. Find the base and the heightif the Island has an area of 64m^2 b = h/2 A = (1/2)b*h 64 = (1/2)(h/2)h 64 = (h/2)^2 h/2 = 8 h = 16 meters